# Schrödinger symmetry: a historical review

### v1

@@ 8 Conformal Galilean Algebra
which is a massless Klein-Gordon equation in $d+2$ dimensions (and not a massive Klein-Gordon equation in $d+1$ time-space dimensions). In the new coordinates $\xi_{-1}=\frac{u}{c},\xi_{0}=ct$ and $\xi_{j}=r_{j}$ with $j=1,\ldots,d$ and $\Psi(\bm{\xi})=\chi(u,t,\bm{r})$, eq. (8.85) becomes $\partial_{\mu}\partial^{\mu}\Psi(\bm{\xi})=0$. Its dynamical symmetry is the conformal group in its usual form, with generators
@@
Figure 2: Left: Root diagramme of the complex Lie algebra $B_{2}$, with the generators $H,P,M,D,B,K$ of table 2 and the four additional ones $V_{\pm},W,N$. Right: the three minimal standard parabolic sub-algebras $\widetilde{\mathfrak{sch}}(1)=\mathfrak{sch}(1)\oplus\mathbb{C}N$, $\widetilde{\mathfrak{age}}(1)=\mathfrak{age}(1)\oplus\mathbb{C}N$ and $\widetilde{\mbox{\sc cga}}(1)=\mbox{\sc cga}(1)\oplus\mathbb{C}N$.
In order to understand the meaning of these result we give in figure 2, via a root diagramme, an overview of the commutator relations of the conformal algebra in $(1+1+1)$ dimensions, which is isomorphic to the complex Lie algebra $B_{2}$ [128]. This illustrates that each generator $\mathscr{X}\in B_{2}$ can be linked to a root vector $\mathfrak{x}$ [^39] If $\mathfrak{x}_{1}+\mathfrak{x}_{2}=\mathfrak{x}_{3}$, then $\bigl{[}\mathscr{X}_{1},\mathscr{X}_{2}\bigr{]}=\mathscr{X}_{3}$, up to a constant factor and up to a linear combination of the roots of the Cartan sub-algebra $\mathfrak{h}=\bigl{\langle}D,N\bigr{\rangle}$. But if $\mathfrak{x}_{1}+\mathfrak{x}_{2}$ falls outside the root diagramme, then $\bigl{[}\mathscr{X}_{1},\mathscr{X}_{2}\bigr{]}=0$. Each convex set leads to a Lie sub-algebra. Lie algebras with different root diagrammes, up to Weyl transformations, are not isomorphic [128]. For clarity, we repeat on the right of figure 2 the root diagramme for the Schrödinger algebra $\widetilde{\mathfrak{sch}}(1)=\mathfrak{sch}(1)\oplus\mathbb{C}N$. Now, the contraction procedure (8.89) leads to a different algebra, called $\mbox{\sc cga}(1)$, whose root diagramme is also indicated on the right of figure 2. Therefore, we have in the non-relativistic limit a projection [93]
@@
This should be compared with the algebras of time-space transformations constructed in section 7. There, it was seen that both the Schrödinger algebra $\mathfrak{sch}(d)$ and the conformal galilean algebra $\mbox{\sc cga}(d)$ may arise either from a study of possible time-space transformations respecting scale-invariance or else from the admissible form of geodesic curves. We now see that these two algebras are also the two main parabolic sub-algebras of the complex conformal Lie algebra $B_{2}$. [^41] Their common sub-algebra $\mathfrak{age}(d)$ was thought to be related to physical ageing, because of the absence of the time-translation generator $H$ [93]. Section 10 deals with applications of Schrödinger-invariance to physical ageing.
The relationship of the conformal galilean algebra with either ortho- or meta-conformal algebras may be illustrated in yet a different way. In $(1+1)$ time-space dimensions (using a more systematic notation of generators [^42] In $(1+1)$ time-space dimensions the isomorphism of ortho- and meta-conformal algebras can be seen as follows [25]. In complex light-cone coordinates $z=t+{\rm i}\mu r$, let $\ell_{n}=-z^{n+1}\partial_{z}+\Delta z^{n}$ and similarly for ¯$\bar{\ell}_{n}$, where $\Delta,\overline{\Delta}$ are the conformal weights. The ortho-conformal generators are $X_{n}=\ell_{n}+\bar{\ell}_{n}$ and $Y_{n}={\rm i}\mu\bigl{(}\ell_{n}-\bar{\ell}_{n}\bigr{)}$. The meta-conformal generators are $X_{n}=\ell_{n}+\bar{\ell}_{n}$ and $Y_{n}=\mu\ell_{n}$. in analogy with table 2 for the Schrödinger algebra) one has for the ortho- and meta-conformal algebras, respectively, the commutators (with $n,n^{\prime}\in\mathbb{Z}$)
@@
where the constraints $x_{1}=x_{2}$ and $\gamma_{1}=\gamma_{2}$ hold. [^43] In the $\mu\to 0$ limit, both ortho- and meta-conformal forms (8.95) reduce to the conformal galilean correlator [83]. For ortho-conformal invariance, the conformal weight $\Delta=\bigl{(}x_{1}-{\rm i}\gamma_{1}/\mu\bigr{)}/2$. Three-point functions $\bigl{\langle}\phi_{1}(t_{1},\bm{r}_{1})\phi_{2}(t_{2},\bm{r}_{2})\phi_{3}(t_{3},\bm{r}_{3})\bigr{\rangle}$ can be fixed similarly [121]. In contrast with the Schrödinger algebras, which predicts co-variant response functions, the co-variant $n$-point functions found from these three algebras are correlation functions [129]. The qualitative behaviour of the associated scaling functions is shown in figure 3a. For large arguments of the scaling variable $u$, both ortho- and meta-conformal correlators decay algebraically, whereas the meta-conformal correlator has an exponential decay. On the other hand, for $u$ small, both meta-conformal and conformal galilean correlators have a pic and are not differentiable at $u=0$, whereas the ortho-conformal correlator has a rounded form. [^44] For $(1+2)D$ meta-conformal invariance, the correlator interpolates between the $(1+1)D$ meta-conformal and the ortho-conformal correlator [83]. In figure 3b, the scaling functions are compared with the one obtained from Schrödinger-invariance (a response function) which clearly highlights their difference ($f(u)\sim e^{-u^{2}}$ has a gaussian form for Schrödinger-invariance).
@@ 9 Super-selection rules
We now revisit the conservation of mass. It is not simply some dynamical symmetry but has deep connections with central extensions of the time-space symmetry algebra.
@@
since the velocity $\bm{v}$ is arbitrary. Therefore the total mass $M$ of a non-relativistic system is always kept fixed, which is obtained here from the non-centrally extended representation (9.98). This mass conservation was established by Lavoisier more than 200 years ago, stating that [103]
@@
2. This become very different when one goes over to non-relativistic quantum mechanics. For a free particle, the entire information is a contained in the wave equation $\psi(t,x)$ which obeys the wave equation (for notational simplicity in $d=1$ space dimensions)
@@
$BC^{[n]}$ $=$ $\left[-(t_{1}-t_{n})\partial_{x_{1}}-\ldots-(t_{n-1}-t_{n})\partial_{x_{n-1}}-\frac{{\rm i}}{\hbar}\bigl{(}m_{1}(x_{1}-x_{n})+\ldots+m_{n-1}(x_{n-1}-x_{n})\bigr{)}\right.$
@@
We compare the root diagrammes in figure 2. The standard galilean algebra is spanned by $\langle B,P,M\rangle\stackrel{{\scriptstyle\wedge}}{{=}}\langle Y_{\frac{1}{2}},Y_{-\frac{1}{2}},M\rangle$, see eq. (9.108), where the central extension $M$ was already included. But the space-transformations of the conformal galilean algebra are spanned by $\langle K,B,M\rangle\stackrel{{\scriptstyle\wedge}}{{=}}\langle Y_{1},Y_{0},Y_{-1}\rangle$ as given by (9.110). Therefore it is clear from figure 2 (or eq. (9.110)) that $\bigl{[}Y_{m},Y_{m^{\prime}}\bigr{]}=0$ and no central extension exists in this case. For the two-point function $C^{[2]}$, it can be easily shown, from the co-variance conditions $Y_{m}C^{[2]}=0$, that the two ‘rapidities’ are equal: $\gamma_{1}=\gamma_{2}$ [121, 136, 137]. Hence the physical rôle of the ‘masses’ $M_{j}$ and the ‘rapidities’ $\gamma_{j}$ is different. [^47] For the infinite-dimensional extension of the conformal galilean algebra, the usual central extensions of Virasoro form are of course admissible. See figure 3 for the comparison of the forms of the two-point scaling functions, according to ortho-conformal, meta-conformal, conformal galilean and Schrödinger invariance.
@@ 10 Physical ageing
Consider a many-body system whose equilibrium state is either critical (with dynamically created long-range correlations) or else has more than one distinct but equivalent equilibrium states. Roughly speaking, physical ageing arises when the time-evolution starts from an initial state which is different from the equilibrium state. For example, one might obtain this situation via quenching a system from a fully disordered initial state to a state either onto or else below a critical temperature $T_{c}>0$, see figure 4. After the quench, the system is far from equilibrium, since it is no longer at a stable minimum of the free energy. Ageing can be monitored through the correlations of the time-space-dependent order-parameter $\phi(t,\bm{r})$. One measures for instance the single-time correlator or the two-time auto-correlator
@@
where $\Delta_{\bm{r}}$ is the spatial laplacian and fluctuations enter through the centred gaussian white noise with variance $\left\langle\eta(t,\bm{r})\eta(t^{\prime},\bm{r}^{\prime})\right\rangle=2T{\cal M}\delta(t-t^{\prime})\delta(\bm{r}-\bm{r}^{\prime})$. Following [144], we shall use this simple model, with its linear Langevin equation, to illustrate some of the main aspects of dynamical Schrödinger symmetry. Clearly, the noise term in (10.113) breaks any time-space symmetry beyond simple translation- and rotation-invariance. Hence the noisy eq. (10.113), as it stands, cannot be Schrödinger-invariant.
@@
We note an important feature of Schrödinger-invariance: the requirement of co-variance fixes directly response functions, such as $R=\bigl{\langle}h\widetilde{h}\bigr{\rangle}$, because they are compatible with the Bargmann super-selection rule (10.116). On the other hand, a co-variance requirement imposed on a correlation function, such as $C=\bigl{\langle}hh\bigr{\rangle}$, would force it to vanish, because the Bargmann super-selection rule (10.116) cannot be satisfied. Correlators will always be obtained by reducing them to higher response functions [145]. The causality of response functions can be systematically derived from a detailed analysis of the time-space representations [93, 129].
@@
This reproduces simulations in many models of phase-ordering, see [121].
2. For a critical quench to $T=T_{c}$, in general the dynamic exponent $z\neq 2$. Since the form of the auto-response functions, determined from co-variance, only depends on $\lambda/z$, that part of the theory can still be used, to a good degree of precision [121, 150]. However, there are indications that a better choice of representation might be a logarithmic one – in analogy to logarithmic conformal field theory [151, 152], where the scaling operators become at least two-component vectors and the scaling dimensions $x$ are replaced by Jordan matrices. Such logarithmic representations have been constructed for the Schrödinger algebra [^52] Analogous constructions also exist for the conformal galilean algebra, including its ‘exotic’ central extension [155, 156, 157]. [153, 154] and indeed permit a much improved agreement with simulational data of response functions in several critical models ($1D$ critical directed percolation [158], the $1D/2D$ Kardar-Parisi-Zhang equation [159, 160] and the $2D$ critical Ising model [161]).
@@ 11 Conclusions
Acknowledgements: This project was initiated jointly with Christian Duval before his untimely death. We would like to thank Gary Gibbons for his interest, advice, and his contribution at the early stages of this work. MH was supported by the french ANR UNIOPEN (ANR-22-CE30-0004-01). PMZ was partially supported by the National Natural Science Foundation of China (Grant No. 11975320).

### v3

@@ 8 Conformal Galilean Algebra
which is a massless Klein-Gordon equation in $d+2$ dimensions (and not a massive Klein-Gordon equation in $d+1$ space-time dimensions). In the new coordinates $\xi_{-1}=\frac{u}{c},\xi_{0}=ct$ and $\xi_{j}=r_{j}$ with $j=1,\ldots,d$ and $\Psi(\bm{\xi})=\chi(u,t,\bm{r})$, eq. (8.85) becomes $\partial_{\mu}\partial^{\mu}\Psi(\bm{\xi})=0$. Its dynamical symmetry is the conformal group in its usual form, with generators
@@
Figure 2: Left: Root diagram of the complex Lie algebra $B_{2}$, with the generators $H,P,M,D,B,K$ of table 2 and the four additional ones $V_{\pm},W,N$. Right: the three minimal standard parabolic sub-algebras $\widetilde{\mathfrak{sch}}(1)=\mathfrak{sch}(1)\oplus\mathbb{C}N$, $\widetilde{\mathfrak{age}}(1)=\mathfrak{age}(1)\oplus\mathbb{C}N$ and $\widetilde{\mbox{\sc cga}}(1)=\mbox{\sc cga}(1)\oplus\mathbb{C}N$.
In order to understand the meaning of these result we give in figure 2, via a root diagram, an overview of the commutator relations of the conformal algebra in $(1+1+1)$ dimensions, which is isomorphic to the complex Lie algebra $B_{2}$ [136]. This illustrates that each generator $\mathscr{X}\in B_{2}$ can be linked to a root vector $\mathfrak{x}$ [^40] If $\mathfrak{x}_{1}+\mathfrak{x}_{2}=\mathfrak{x}_{3}$, then $\bigl{[}\mathscr{X}_{1},\mathscr{X}_{2}\bigr{]}=\mathscr{X}_{3}$, up to a constant factor and up to a linear combination of the roots of the Cartan sub-algebra $\mathfrak{h}=\bigl{\langle}D,N\bigr{\rangle}$. But if $\mathfrak{x}_{1}+\mathfrak{x}_{2}$ falls outside the root diagram, then $\bigl{[}\mathscr{X}_{1},\mathscr{X}_{2}\bigr{]}=0$. Each convex set leads to a Lie sub-algebra. Lie algebras with different root diagrams, up to Weyl transformations, are not isomorphic [136]. For clarity, we repeat on the right of figure 2 the root diagram for the Schrödinger algebra $\widetilde{\mathfrak{sch}}(1)=\mathfrak{sch}(1)\oplus\mathbb{C}N$. Now, the contraction procedure (8.89) leads to a different algebra, called $\mbox{\sc cga}(1)$, whose root diagram is also indicated on the right of figure 2. Therefore, we have in the non-relativistic limit a projection [101]
@@
This should be compared with the algebras of space-time transformations constructed in section 7. There, it was seen that both the Schrödinger algebra $\mathfrak{sch}(d)$ and the conformal Galilean algebra $\mbox{\sc cga}(d)$ may arise either from a study of possible space-time transformations respecting scale-invariance or else from the admissible form of geodesic curves. We now see that these two algebras are also the two main parabolic sub-algebras of the complex conformal Lie algebra $B_{2}$. [^42] Their common sub-algebra $\mathfrak{age}(d)$ was thought to be related to physical ageing, because of the absence of the time-translation generator $H$ [101]. Section 10 deals with applications of Schrödinger-invariance to physical ageing.
The relationship of the conformal Galilean algebra with either ortho- or meta-conformal algebras may be illustrated in yet a different way. In $(1+1)$ space-time dimensions (using a more systematic notation of generators [^43] In $(1+1)$ space-time dimensions the isomorphism of ortho- and meta-conformal algebras can be seen as follows [33]. In complex light-cone coordinates $z=t+{\rm i}\mu r$, let $\ell_{n}=-z^{n+1}\partial_{z}+\Delta z^{n}$ and similarly for ¯$\bar{\ell}_{n}$, where $\Delta,\overline{\Delta}$ are the conformal weights. The ortho-conformal generators are $X_{n}=\ell_{n}+\bar{\ell}_{n}$ and $Y_{n}={\rm i}\mu\bigl{(}\ell_{n}-\bar{\ell}_{n}\bigr{)}$. The meta-conformal generators are $X_{n}=\ell_{n}+\bar{\ell}_{n}$ and $Y_{n}=\mu\ell_{n}$. in analogy with table 2 for the Schrödinger algebra) one has for the ortho- and meta-conformal algebras, respectively, the commutators (with $n,n^{\prime}\in\mathbb{Z}$)
@@
where the constraints $x_{1}=x_{2}$ and $\gamma_{1}=\gamma_{2}$ hold. [^44] In the $\mu\to 0$ limit, both ortho- and meta-conformal forms (8.95) reduce to the conformal Galilean correlator [91]. For ortho-conformal invariance, the conformal weight $\Delta=\bigl{(}x_{1}-{\rm i}\gamma_{1}/\mu\bigr{)}/2$. Three-point functions $\bigl{\langle}\phi_{1}(t_{1},\bm{r}_{1})\phi_{2}(t_{2},\bm{r}_{2})\phi_{3}(t_{3},\bm{r}_{3})\bigr{\rangle}$ can be fixed similarly [129]. In contrast with the Schrödinger algebras, which predicts co-variant response functions, the co-variant $n$-point functions found from these three algebras are correlation functions [137]. The qualitative behaviour of the associated scaling functions is shown in figure 3a. For large arguments of the scaling variable $u$, both ortho- and meta-conformal correlators decay algebraically, whereas the meta-conformal correlator has an exponential decay. On the other hand, for $u$ small, both meta-conformal and conformal Galilean correlators are not differentiable at $u=0$, whereas the ortho-conformal correlator has a rounded form. [^45] For $(1+2)D$ meta-conformal invariance, the correlator interpolates between the $(1+1)D$ meta-conformal and the ortho-conformal correlator [91]. In figure 3b, the scaling functions are compared with the one obtained from Schrödinger-invariance (a response function) which clearly highlights their difference ($f(u)\sim e^{-u^{2}}$ has a Gaussian form for Schrödinger-invariance).
@@ 9 Super-selection rules
We now revisit the conservation of mass. It is not simply some dynamical symmetry but has deep connections with central extensions of the space-time symmetry algebra.
@@
since the velocity $\bm{v}$ is arbitrary. Therefore the total mass $M$ of a non-relativistic system is always kept fixed, which is obtained here from the non-centrally extended representation (9.98). This mass conservation was established by Lavoisier more than 200 years ago, stating that [^46] Nothing is created, neither artificially, nor in Nature, and one may pose as a principle that in all operations there is the same quantity of matter before and after the operation. [111]
@@
2. This becomes very different when one goes over to non-relativistic quantum mechanics. For a free particle, the entire information is contained in the wave equation $\psi(t,x)$ which obeys the wave equation (for notational simplicity in $d=1$ space dimensions)
@@
$BC^{[n]}$ $=$ $\left[-(t_{1}-t_{n})\partial_{x_{1}}-\ldots-(t_{n-1}-t_{n})\partial_{x_{n-1}}\right.$
$\left.-\frac{{\rm i}}{\hbar}\Bigl{(}m_{1}(x_{1}-x_{n})+\ldots+m_{n-1}(x_{n-1}-x_{n})\Bigr{)}\right.$
@@
We compare the root diagrams in figure 2. The standard Galilean algebra is spanned by $\langle B,P,M\rangle\stackrel{{\scriptstyle\wedge}}{{=}}\langle Y_{\frac{1}{2}},Y_{-\frac{1}{2}},M\rangle$, see eq. (9.108), where the central extension $M$ was already included. But the space-transformations of the conformal Galilean algebra are spanned by $\langle K,B,M\rangle\stackrel{{\scriptstyle\wedge}}{{=}}\langle Y_{1},Y_{0},Y_{-1}\rangle$ as given by (9.110). Therefore it is clear from figure 2 (or eq. (9.110)) that $\bigl{[}Y_{m},Y_{m^{\prime}}\bigr{]}=0$ and no central extension exists in this case. For the two-point function $C^{[2]}$, it can be easily shown, from the co-variance conditions $Y_{m}C^{[2]}=0$, that the two ‘rapidities’ are equal: $\gamma_{1}=\gamma_{2}$ [129, 144, 145]. Hence the physical rôle of the ‘masses’ $M_{j}$ and the ‘rapidities’ $\gamma_{j}$ is different. [^49] For the infinite-dimensional extension of the conformal Galilean algebra, the usual central extensions of Virasoro form are of course admissible. See figure 3 for the comparison of the forms of the two-point scaling functions, according to ortho-conformal, meta-conformal, conformal Galilean and Schrödinger invariance.
@@ 10 Physical ageing
Consider a many-body system whose equilibrium state is either critical (with dynamically created long-range correlations) or else has more than one distinct but equivalent equilibrium states. Roughly speaking, physical ageing arises when the time-evolution starts from an initial state which is different from the equilibrium state. For example, one might obtain this situation via quenching a system from a fully disordered initial state to a state either onto or else below a critical temperature $T_{c}>0$, see figure 4. After the quench, the system is far from equilibrium, since it is no longer at a stable minimum of the free energy. Ageing can be monitored through the correlations of the space-time-dependent order-parameter $\phi(t,\bm{r})$. One measures for instance the single-time correlator or the two-time auto-correlator
@@
where $\Delta_{\bm{r}}$ is the spatial Laplacian and fluctuations enter through the centred Gaussian white noise with variance $\left\langle\eta(t,\bm{r})\eta(t^{\prime},\bm{r}^{\prime})\right\rangle=2T{\cal M}\delta(t-t^{\prime})\delta(\bm{r}-\bm{r}^{\prime})$. Following [152], we shall use this simple model, with its linear Langevin equation, to illustrate some of the main aspects of dynamical Schrödinger symmetry. Clearly, the noise term in (10.113) breaks any space-time symmetry beyond simple translation- and rotation-invariance. Hence the noisy eq. (10.113), as it stands, cannot be Schrödinger-invariant.
@@
We note an important feature of Schrödinger-invariance: the requirement of co-variance fixes directly response functions, such as $R=\bigl{\langle}h\widetilde{h}\bigr{\rangle}$, because they are compatible with the Bargmann super-selection rule (10.116). On the other hand, a co-variance requirement imposed on a correlation function, such as $C=\bigl{\langle}hh\bigr{\rangle}$, would force it to vanish, because the Bargmann super-selection rule (10.116) cannot be satisfied. Correlators will always be obtained by reducing them to higher response functions [153]. The causality of response functions can be systematically derived from a detailed analysis of the space-time representations [101, 137].
@@
This reproduces simulations in many models of phase-ordering, see [129, 158].
2. For a critical quench to $T=T_{c}$, in general the dynamic exponent $z\neq 2$. Since the form of the auto-response functions, determined from co-variance, only depends on $\lambda/z$, that part of the theory can still be used, to a good degree of precision [129, 159]. However, there are indications that a better choice of representation might be a logarithmic one – in analogy to logarithmic conformal field theory [160, 161], where the scaling operators become at least two-component vectors and the scaling dimensions $x$ are replaced by Jordan matrices. Such logarithmic representations have been constructed for the Schrödinger algebra [^54] Analogous constructions also exist for the conformal Galilean algebra, including its ‘exotic’ central extension [164, 165, 166]. [162, 163] and indeed permit a much improved agreement with simulation data of response functions in several critical models ($1D$ critical directed percolation [167], the $1D/2D$ Kardar-Parisi-Zhang equation [168, 169] and the $2D$ critical Ising model [170]).
@@ 11 Conclusions
After this paper was submitted, we were informed of several further researches [187, 188, 189, 190, 191, 192, 193] related to Schrödinger symmetry.
Acknowledgements: This project was initiated jointly with Christian Duval before his untimely death. We would like to thank Gary Gibbons for his interest, advice, and his contribution at the early stages of this work. MH was supported by the French ANR UNIOPEN (ANR-22-CE30-0004-01). PMZ was partially supported by the National Natural Science Foundation of China (Grant No. 12375084).